OpenStudy (anonymous):

5 1/4 divided by 1 5/9 3 1/4 3 3/8 4 1/4 8 1/6 will some one please check my answers??

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

i think its the first one is that right??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

5 1/4 = 5(4) + 1 /4 = 21/4

OpenStudy (anonymous):

1 5/9 = 14/9

OpenStudy (anonymous):

21/4 * 9/14

OpenStudy (anonymous):

3/4 * 9/2

OpenStudy (anonymous):

27/8

OpenStudy (anonymous):

so second one is answer

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

I will show you clearly wait

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Convert both mixed fractions into simple fractions

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay i got 21/4 and 14/9

OpenStudy (the_fizicx99):

5 1/4 divided by 1 5/9 Convert them into simple fractions, then do the Keep, Switch Flip Meaning Keep the first fraction Switch the division sign to multiplication and flip the second fraction 5*4 + 1 & 1*9 + 5 21/4 * 9/14 3/4 * 9/2 27/8 8 goes into 27 3 time and a left over of 3 and keep the denominator So, 3 3/8

OpenStudy (anonymous):

what about this one??? 3 1/2 times (- 4 4/5) -25 3/5 -20 4/15 20 4/15 25 3/5

OpenStudy (the_fizicx99):

8 goes into 27; 3 times (8*3 = 24) and a left over of 3. 24 + 3 = 27

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

3 1/2 => 7/2 -4 4/5 = > -24/5 7/2(-24/5) -84/5

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I dont see answer for that in Options ;)

OpenStudy (anonymous):

I think I did it wrong

OpenStudy (the_fizicx99):

Convert to simple form Then just multiply across. 3 * 2 + 1 and -4 * 5 + 4 6 + 1 = 7 and -24/5 Keep the same denominator, 7/2 * -24/5 -84/5 -16 4/5 <-- is what I got, not one of the options though..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks